# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v1

@@ 4 Method: Sampling Then Accepting
In this section, we propose the Sampling One Then Accepting (STA-1) method, and discuss detecting the STA-1 generated text using the $z$-test. Theoretically, we show that STA-1 is unbiased and its type II error of the $z$-test has statistical guarantees. Next, we analyze all unbiased watermarks under the low-entropy protocol in Section 3. We finally introduce Sampling M Then Accepting (STA-M), an extension of STA-1.
@@ 4.1 Sampling One Then Accepting
We start by proposing the Sampling One Then Accepting (STA-1) method in Algorithm 1, which is always unbiased and easy to analyze. First, the hash value of the last generated token is computed and employed as the seed of a random number generator (RNG). We use the RNG to divide the token set into a green and a red list [12]. Next, we sample from the original LLM output distribution (as depicted in Line 4 of Algorithm 1), accept the sampling if the token is in the green list (as depicted in Line 5 and Line 6 in Algorithm 1), sample again if the token is in the red list (as depicted in Line 7 in Algorithm 1), and the second sampling is always accepted.
@@
Input: A pretrained LLM $P_{M}$, a key $k\in K$, the proportion of green list $\gamma\in(0,1)$, and a prompt $x^{-N_{p}:0}$
@@
STA-1 is a simple but effective watermarking method. The properties of STA-1 include: (1) STA-1 is an unbiased watermark; (2) The number of green list tokens in STA-1 generated texts has a lower bound on its mean and an upper bound on its variance, which provides statistical guarantees on the detection of STA-1; (3) STA-1 has a lower risk for low-entropy generation compared to previous work. In deriving theoretical results, we assume that the key $k$ is randomly sampled from a uniform distribution. Therefore, the random partition of green and red lists associated with this key is also uniform [12, 11]. We start by analyzing the unbiased characteristic of STA-1.
@@ 4.1.1 Statistical test guarantees of STA-1
Detecting the STA-1 generated text. The detection of STA-1 compares the empirical proportion of green list tokens in the given text against the green list proportion $\gamma$ [12]. We employ the $z$-test where the null hypothesis ($H_{0}$) is that the text is generated without knowing the green-red list partition [12]. Denote $|S|_{G}$ as the number of green list tokens in this text. Under $H_{0}$, $|S|_{G}$ follows a Bernoulli distribution $B(T,\gamma)$ with a mean of $\gamma T$ and a variance of $\gamma(1-\gamma)T$. The $z$-score is calculated with the empirical $|S|_{G}$ as
@@
Remark 2. Compared to the Spike entropy proposed by Kirchenbauer et al. [12], the Gini index is a commonly used metric in machine learning to measure the uncertainty of a probability distribution, such as CART decision tree [3].
@@ 4.1.2 Discussion on the low-entropy protocol
Previous work claims that unbiased watermarks can avoid the tradeoff between watermark strength and text quality [11]. We challenge this claim by first considering the following example.
Example 2. Assuming that the token set only includes two tokens $\mathcal{V}=\{A,B\}$, at a typical step, an LLM outputs the probability of generating $A$ ($p_{A}$) and $B$ ($p_{B}$) as $(p_{A},p_{B})=(0.8,0.2)$. Consider the following two unbiased watermarks. (1) With a probability of 0.2 always generating $B$ and with a probability of 0.8 always generating $A$. (2) With a probability of 0.5 probabilities becoming $(p_{A},p_{B})=(0.9,0.1)$ and with the other probability of 0.5 becoming $(p_{A},p_{B})=(0.7,0.3)$.
We strongly recommend readers who are unfamiliar with risk-averse or utility theory to Appendix C for a conventional example in finance. In Example 2, one can view the prompt as ‘The quick brown fox jumps over the lazy’, $A$ as the token ‘dog’, and $B$ as all other tokens. It is easy to show that watermarks (1) and (2) are both unbiased. However, risk-averse people [22] will prefer watermark (2) because watermark (2) does not have a probability that only $B$ will be sampled. $B$ represents unsatisfactory outputs in low-entropy scenarios which could significantly harm text quality, and we want the risk of sampling $B$ as low as possible. Again, we refer readers to Appendix C for a better understanding of the analysis. At any generation step, let $x_{max}$ denote the token with the maximum probability $p_{max}$. We measure the risk by the variance [27] of $p_{max}^{w,k}$ among watermark keys, where $p_{max}^{w,k}$ denotes the altered value of $p_{max}$ with a watermarking method and a key $k$. We show that STA-1 has a lower risk compared to previous unbiased watermarks in the following theorem.
Theorem 3. Assume $11-\alpha\leq p_{max}<$ 1, where $\alpha$ represents the partition hyperparameter in Dipmark. For the low-entropy protocol in Section 3, the STA-1 method has a lower variance in the probability of generating $x_{max}$ compared to other unbiased methods (including Dipmark and $\gamma$-reweight) [11, 33]. Formally,
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right],$ (9)
@@
Proof. See Appendix B.3.
Example 3. We show a numerical example with $p_{max}=0.8$. For STA-1, based on the proof of the theorem, if the proportion of the green list is 0.5, $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=6.4\times 10^{-3}$. For Dipmark ($\alpha\in[0.2,0.5]$) and $\gamma$-reweight, the variance is $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]=\frac{1}{75}\approx 0.013$.
@@ 4.2 Sampling M Then Accepting
A low-entropy scenario indicates a low Gini index which weakens the watermark strength based on Theorem 2. To enhance the watermark strength, we propose the Sampling M Then Accepting (STA-M) method, an extension of STA-1. STA-M employs a heuristic threshold $\tau$ for entropy at each generation step. In detail, at generation step $t$, we first calculate the entropy $\tau^{t}$ of the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. If it shows low entropy $\tau^{t}\leq\tau$, we apply STA-1 at this generation step; if it shows high entropy $\tau^{t}>\tau$, we repeat sampling if the previously sampled token is in the red list, and the procedure repeats at most $M$ times. The detailed algorithm and analysis of STA-M can be found in Appendix D.
@@ 5.1 Experimental Setup
Datasets and metrics. We employed two public datasets: C4 subset [24, 12] for news-like text generation and HumanEval [4] for code generation. We evaluated the performance of different watermarking methods on text quality and watermark strength. For watermark strength, we set the $z$ threshold as 2 and 2.5 and report the F1-score and AUC of watermark detection. For text quality, we measured perplexity (PPL) and coherence [9] for generations on C4; We computed PPL and pass@$k$ scores of code generations [4] for HumanEval. We refer readers to Appendix E.1 for more dataset details and the prompt used in datasets.
Baselines. We chose KGW as the biased watermark baseline [12], $\gamma$-reweight [11] and Dipmark [33] as the unbiased watermark baselines. Specifically, we set KGW with a fixed green list proportion $\gamma=0.5$ and diverse logit increments $\delta\in\{1,1.5,2\}$. The partition parameter of Dipmark was set as $\alpha\in\{0.3,0.4,0.5\}$. When $\alpha=0.5$, we report this result as $\gamma$-reweight. Note that $\gamma$-reweight [11] does not include a $z$-test. We implemented the $z$-score in Dipmark [33] by counting the number of tokens in the latter portion of the token set. We also show results without watermarking techniques.
Implementation details. We utilized different variants of LLaMA-2-7B [29] as our generative models, and LLaMA-2-13B to compute perplexity. For hyperparameters in STA-M, we set $M\in\{4,8,16\}$ and two entropy thresholds $\tau$ for different datasets. We conducted a robustness check on $\tau$ in Appendix E.2 and selected different $\tau$s for different datasets in the final experiment. We refer readers to Appendix E.1 for more details on implementation.
Table 1: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the C4 Dataset.
Text Quality Watermark Strength
$z=2.0$ $z=2.5$
@@
No Watermark 7.409 0.605 0.049 0.495 0.012 0.499
KGW($\delta$=1) 7.617 0.602 0.956 0.957 0.928 0.932
KGW($\delta$=1.5) 7.870 0.600 0.982 0.982 0.990 0.990
KGW($\delta$=2) 8.221 0.602 0.980 0.980 0.993 0.993
Dipmark($\alpha$=0.3) 7.451 0.604 0.923 0.924 0.893 0.901
Dipmark($\alpha$=0.4) 7.451 0.604 0.951 0.951 0.943 0.945
$\gamma$-reweight 7.437 0.611 0.969 0.969 0.964 0.965
STA-1 7.343 0.600 0.957 0.956 0.961 0.961
STA-4($\tau$=1.35) 7.735 0.588 0.981 0.981 0.991 0.991
STA-8($\tau$=1.35) 8.161 0.589 0.969 0.968 0.991 0.991
STA-16($\tau$=1.35) 7.988 0.585 0.974 0.974 0.992 0.992
5.2 Results on C4
For the C4 dataset, each method generates at least 500 text sequences with at least $200\pm 5$ tokens [12]. Table 1 demonstrates the text quality and watermark strength for each method, and we present generated text examples in Appendix E.3. As depicted in Table 1, the proposed STA-1 method achieves comparable perplexity and coherence compared to no watermark generation. The text quality results are consistent with other unbiased watermarks including Dipmark and $\gamma$-reweight, showing STA-1 is also unbiased empirically. In terms of watermark strength, STA-M ($M\in\{4,8,16\}$) outperforms unbiased watermarks including STA-1 and Dipmark, and has similar watermark strength as biased watermarks KGW ($\delta\in\{1.5,2\}$). Overall in the high-entropy generation task, the unbiased STA-1 method is comparable to other unbiased watermarks; The STA-M method can improve the watermark strength by sacrificing minor text quality.
Table 2: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the HumanEval Dataset.
@@ 5.2 Results on C4
No Watermark 3.275 0.147 0.405 0.537 0.114 0.494 0.072 0.497
KGW($\delta$=1) 3.288 0.119 0.326 0.415 0.471 0.643 0.416 0.627
KGW($\delta$=1.5) 3.285 0.104 0.308 0.427 0.720 0.770 0.650 0.730
KGW($\delta$=2) 3.650 0.077 0.254 0.372 0.757 0.795 0.733 0.785
Dipmark($\alpha$=0.3) 3.081 0.142 0.392 0.512 0.518 0.665 0.423 0.625
Dipmark($\alpha$=0.4) 3.020 0.135 0.393 0.512 0.516 0.668 0.429 0.634
$\gamma$-reweight 3.314 0.130 0.371 0.488 0.522 0.671 0.479 0.655
STA-1 3.006 0.155 0.394 0.494 0.526 0.633 0.442 0.611
STA-4($\tau$=1.95) 3.175 0.151 0.392 0.500 0.633 0.685 0.594 0.679
STA-8($\tau$=1.95) 2.842 0.148 0.399 0.537 0.652 0.703 0.587 0.675
STA-16($\tau$=1.95) 3.024 0.145 0.382 0.476 0.725 0.764 0.640 0.717

### v2

@@ 4 Method: Sampling Then Accepting
In this section, we propose the Sampling One Then Accepting (STA-1) method, and discuss detecting the STA-1 generated text using the $z$-test. Theoretically, we show that STA-1 is unbiased and its type II error of the $z$-test has statistical guarantees. Next, we analyze previous unbiased watermarks and STA-1 under the low-entropy protocol in Section 3. We finally introduce Sampling M Then Accepting (STA-M), an extension of STA-1.
@@ 4.1 Sampling One Then Accepting
We start by proposing the Sampling One Then Accepting (STA-1) method in Algorithm 1, which is always unbiased and easy to analyze. First, the hash value of the last generated token is computed and employed as the seed of a random number generator (RNG). We use the RNG to divide the token set into a green and a red list (Kirchenbauer et al., 2023a). Next, we sample from the original LLM output distribution (as depicted in Line 4 of Algorithm 1), accept the sampling if the token is in the green list (as depicted in Line 5 and Line 6 in Algorithm 1), sample again if the token is in the red list (as depicted in Line 7 and Line 8 in Algorithm 1), and the second sampling is always accepted.
@@
Input: A pretrained LLM $P_{M}$, a watermark key $k\in K$, the proportion of the green list $\gamma\in(0,1)$, and a prompt $x^{-N_{p}:0}$
@@
STA-1 is a simple but effective watermarking method. The properties of STA-1 include: (1) STA-1 is an unbiased watermark; (2) The number of green list tokens in STA-1 generated texts has a lower bound on its mean and an upper bound on its variance, which further provides explicit statistical guarantees for the type II error in the STA-1 detection test; (3) STA-1 has a lower risk for low-entropy generation compared to previous work. In deriving theoretical results, we assume that the key $k$ is randomly sampled from a uniform distribution. Therefore, the random partition of green and red lists associated with this key is also uniform (Kirchenbauer et al., 2023a). We start by analyzing the unbiased characteristic of STA-1.
@@ 4.1.1 Statistical Test Guarantees of STA-1
Detecting the STA-1 generated text. The detection of STA-1 compares the empirical proportion of green list tokens in the given text against the green list proportion $\gamma$ (Kirchenbauer et al., 2023a). We employ the $z$-test where the null hypothesis ($H_{0}$) is that the text is generated without knowing the green-red list partition. Denote $|S|_{G}$ as the number of green list tokens in this text. Under $H_{0}$, $|S|_{G}$ follows a Bernoulli distribution $B(T,\gamma)$ with a mean of $\gamma T$ and a variance of $\gamma(1-\gamma)T$. The $z$-score is calculated with the empirical $|S|_{G}$ as
@@
Remark 2. Compared to the Spike entropy proposed by Kirchenbauer et al. (2023a), the Gini index is a commonly used metric in machine learning to measure the uncertainty of a probability distribution, such as CART decision tree (Breiman, 2017).
@@
Having established the mean and variance bounds for $|S|_{G}$, with an additional condition, we derive from Theorem 2 a corollary that provides an explicit upper bound on the type II error of the $z$-test in detecting STA-1.
Corollary 1. Given that Theorem 2 holds, if $Gini^{*}>\tilde{z}/\sqrt{\gamma(1-\gamma)T}$, we have the type II error
$P\left(\frac{|S|_{G}-\gamma T}{\sqrt{\gamma(1-\gamma)T}}\leq\tilde{z}\bigg{|}H_{a}\right)\leq\frac{\overline{\mathbb{V}}}{\overline{\mathbb{V}}+(\underline{\mathbb{E}}-\gamma T-\tilde{z}\sqrt{\gamma(1-\gamma)T})^{2}},$ (8)
where $\tilde{z}$ is the $z$ threshold value, $\underline{\mathbb{E}}$ and $\overline{\mathbb{V}}$ are the lower bound and upper bound values on $\mathbb{E}(|S|_{G})$ and $\mathbb{V}(|S|_{G})$ as established in Theorem 2, respectively.
Proof. See Appendix C.3.
The additional condition requires that $Gini^{*}$ must not be excessively low given the threshold value $\tilde{z}$. A higher $Gini^{*}$ increases $\underline{\mathbb{E}}$ and decreases $\overline{\mathbb{V}}$, resulting in a reduced upper bound on the type II error. Therefore, the test has higher statistical power in high-entropy scenarios.
@@ 4.1.2 Discussion on the Low-entropy Protocol
Previous work claims that unbiased watermarks can avoid the tradeoff between watermark strength and text quality (Hu et al., 2024). We challenge this claim by first considering the following example.
Example 2. Assuming that the token set only includes two tokens $\mathcal{V}=\{A,B\}$, at a typical step, an LLM outputs the probability of generating $A$ ($p_{A}$) and $B$ ($p_{B}$) as $(p_{A},p_{B})=(0.8,0.2)$. Consider the following two unbiased watermarks. $W_{1}$: with a probability of 0.2 always generating $B$ and with a probability of 0.8 always generating $A$; $W_{2}$: with a probability of 0.5, the probability distribution becomes $(p_{A},p_{B})=(0.9,0.1)$ and with the other probability of 0.5, becomes $(p_{A},p_{B})=(0.7,0.3)$.
In Example 2, one can view the prompt as ‘The quick brown fox jumps over the lazy’, $A$ as the token ‘dog’, and $B$ as all other tokens. It is easy to show that watermarks $W_{1}$ and $W_{2}$ are both unbiased. However, risk-averse people (Pratt, 1978) will prefer watermark $W_{2}$ because $W_{2}$ does not have a possibility that only $B$ will be sampled. $B$ represents unsatisfactory outputs in low-entropy scenarios which could significantly harm text quality, and we want the risk of sampling $B$ to be as low as possible. We refer readers to Appendix D for a conventional example in finance and a better understanding of the analysis via utility theory. At any generation step, let $x_{max}$ denote the token with the maximum probability $p_{max}$. We measure the risk by the variance (Sharpe, 1998) of $p_{max}^{w,k}$ among watermark keys, where $p_{max}^{w,k}$ denotes the altered value of $p_{max}$ with a watermarking method and a key $k$. We show that STA-1 has a lower risk compared to previous unbiased watermarks in the following theorem.
Theorem 3. Assume $11-\alpha\leq p_{max}<$ 1, where $\alpha$ represents the partition hyperparameter used in Dipmark. For the low-entropy protocol in Section 3, the STA-1 method has a lower variance in the probability of generating $x_{max}$ compared to other unbiased methods (including Dipmark, $\gamma$-reweight, and RDW) (Hu et al., 2024; Wu et al., 2024; Kuditipudi et al., 2023). Formally,
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$ (9)
@@
Proof. See Appendix C.4.
Example 3. We show a numerical example with $p_{max}=0.8$. For STA-1, based on the proof of the theorem, if the proportion of the green list is 0.5, $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=0.0064$. For Dipmark ($\alpha\in[0.2,0.5]$) and $\gamma$-reweight, the variance is $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]=\frac{1}{75}\approx 0.013$. For RDW, the variance is $\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right]=0.16$.
@@ 4.2 Sampling M Then Accepting
A low-entropy scenario indicates a low Gini index which weakens the watermark strength based on Theorem 2. To enhance the watermark strength, we propose the Sampling M Then Accepting (STA-M) method, an extension of STA-1. STA-M employs a heuristic threshold $\tau$ for entropy at each generation step. In detail, at generation step $t$, we first calculate the entropy $\tau^{t}$ of the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. If it shows low entropy $\tau^{t}\leq\tau$, we apply STA-1 at this generation step; if it shows high entropy $\tau^{t}>\tau$, we repeat sampling if the previously sampled token is in the red list, and the procedure repeats at most $M$ times.
The detailed algorithm and analysis of STA-M can be found in Appendix E. According to Remark 3 in Appendix E, STA-M is biased. In low-entropy steps where probabilities are concentrated on a few tokens, actively using STA-M by repeated sampling can skew these probabilities, thereby reducing text quality. On the contrary, in high-entropy steps, since there are more acceptable tokens, the impact of repeated sampling on text quality is weakened. Therefore, STA-M only repeats sampling in high-entropy steps, which could increase watermark strength and largely maintain text quality.
@@ 5.1 Experimental Setup
Datasets and metrics. We employed two public datasets: C4 subset (Raffel et al., 2020; Kirchenbauer et al., 2023a) for news-like text generation and HumanEval (Chen et al., 2021) for code generation. We evaluated the performance of different watermarking methods on text quality and watermark strength. For watermark strength, we set the $z$ threshold as 2 and 2.5 and report the F1-score and AUC of watermark detection. For text quality, we measured perplexity (PPL) and coherence (Gao et al., 2021) for generations on C4; We computed PPL and pass@$k$ scores of code generations (Chen et al., 2021) for HumanEval. We refer readers to Appendix F.1 for more dataset details and the prompt used in each dataset.
Baselines. We chose KGW as the biased watermark baseline (Kirchenbauer et al., 2023a), RDW (Kuditipudi et al., 2023), $\gamma$-reweight (Hu et al., 2024), and Dipmark (Wu et al., 2024) as the unbiased watermark baselines. Specifically, we set KGW with a fixed green list proportion $\gamma=0.5$ and diverse logit increments $\delta\in\{1,1.5,2\}$. We set the watermark key length as 256 in RDW. The partition parameter of Dipmark was set as $\alpha\in\{0.3,0.4,0.5\}$. When $\alpha=0.5$, we report this result as $\gamma$-reweight. Note that $\gamma$-reweight (Hu et al., 2024) does not include a $z$-test. Therefore, we implemented the $z$-score in Dipmark (Wu et al., 2024) for $\gamma$-reweight by counting the number of tokens in the latter portion of the token set. We also show results without watermarking techniques. Also, RDW only contains a permutation test that reports p-values. We set p-value thresholds at 0.05 and 0.01 to approximate two $z$-tests.
Implementation details. We utilized different variants of LLaMA-2-7B (Touvron et al., 2023) as our generative models, and LLaMA-2-13B to compute perplexity. For hyperparameters in STA-M, we set $M\in\{4,8,16\}$ and two entropy thresholds $\tau$ for different datasets. We conducted a robustness check on $\tau$ in Appendix F.2 and selected different $\tau$s for different datasets in the final experiment. For each method, we run 10 times to conduct all pair-wise Tukey tests. Results in the following tables show only average values. We refer readers to Appendix F.1 for more details on implementation.
Table 1: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the C4 Dataset. The best results without statistical differences are shown in bold. The second best results without statistical differences are shown in underline.
Text Quality Watermark Strength Detection
$z=2.0$ $z=2.5$ Time
@@
No Watermark 7.474 0.604 0.046 0.500 0.012 0.500 46s
KGW($\delta$=1) 7.591 0.606 0.961 0.962 0.940 0.944 46s
KGW($\delta$=1.5) 7.844 0.604 0.985 0.984 0.992 0.992 46s
KGW($\delta$=2) 8.091 0.599 0.986 0.986 0.995 0.995 46s
RDW 7.650 0.592 0.982 0.982 0.948 0.950 4h
Dipmark($\alpha$=0.3) 7.415 0.599 0.933 0.935 0.909 0.915 44s
Dipmark($\alpha$=0.4) 7.384 0.601 0.957 0.957 0.954 0.955 44s
$\gamma$-reweight 7.436 0.599 0.961 0.961 0.963 0.963 44s
STA-1 7.387 0.600 0.962 0.961 0.963 0.963 46s
STA-4($\tau$=1.35) 7.711 0.595 0.973 0.972 0.988 0.988 46s
STA-8($\tau$=1.35) 8.006 0.592 0.975 0.975 0.987 0.987 46s
STA-16($\tau$=1.35) 8.199 0.588 0.973 0.972 0.988 0.988 46s
Table 2: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the HumanEval Dataset. The best results without statistical differences are shown in bold. The second best results without statistical differences are shown in underline.
@@
No Watermark 3.041 0.138 0.405 0.537 0.114 0.494 0.072 0.497
KGW($\delta$=1) 3.078 0.135 0.326 0.415 0.471 0.643 0.416 0.627
KGW($\delta$=1.5) 3.499 0.098 0.308 0.427 0.720 0.770 0.650 0.730
KGW($\delta$=2) 3.723 0.098 0.254 0.372 0.757 0.795 0.733 0.785
RDW 3.159 0.134 0.362 0.470 0.408 0.628 0.343 0.604
Dipmark($\alpha$=0.3) 3.037 0.144 0.392 0.512 0.518 0.665 0.423 0.625
Dipmark($\alpha$=0.4) 3.101 0.141 0.393 0.512 0.516 0.668 0.429 0.634
$\gamma$-reweight 3.088 0.142 0.371 0.488 0.522 0.671 0.479 0.655
STA-1 3.006 0.147 0.394 0.494 0.526 0.633 0.442 0.611
STA-4($\tau$=1.95) 3.175 0.135 0.392 0.500 0.633 0.685 0.594 0.679
STA-8($\tau$=1.95) 2.842 0.146 0.399 0.537 0.652 0.703 0.587 0.675
STA-16($\tau$=1.95) 3.024 0.140 0.382 0.476 0.725 0.764 0.640 0.717
5.2 Results on C4
For the C4 dataset, each method generates at least 500 text sequences with at least $200\pm 5$ tokens (Kirchenbauer et al., 2023a). Table 1 demonstrates each method’s text quality, watermark strength, and detection time for 500 generations, and we present generated text examples in Appendix F.3. As depicted in Table 1, the proposed STA-1 method is efficient in detection and achieves comparable perplexity and coherence compared to no watermark generation. The text quality results are consistent with other unbiased watermarks including RDW, Dipmark, and $\gamma$-reweight, showing STA-1 is also unbiased empirically. In terms of watermark strength, STA-M ($M\in\{4,8,16\}$) outperforms all unbiased watermarks and has comparable watermark strength as biased watermarks KGW ($\delta\in\{1.5,2\}$). Overall in the high-entropy generation task, the unbiased STA-1 method is comparable to other unbiased watermarks; The STA-M method can improve the watermark strength by sacrificing minor text quality.